OpenStudy (anonymous):

Which of the following systems is equivalent to the given system? 2/3x - y = 2 x + 1/2y = -3

OpenStudy (advert):

4x - y = 4 (2/3)x + (1/2)y = -1 , multiply by 6: 4x + 3y = -6 Option (c)

OpenStudy (anonymous):

These are the options. 2x - 3y = 6 and 2x + y = 6 6x - 3y = 6 and 2x + y = 10 2x - 3y = 6 and 2x + y = -6

OpenStudy (advert):

hope i helped